At noon mr.Pryme decides to live kumasi for cape coast driving at 50km per hour.at 1:00pm Miss Anna live cape coast for kumasi driving at 60km per hour along the same route.If the two towns are 226km apart at what time will mr.Pryme meet Miss anna.
Found 2 solutions by jorel555, MathTherapy:
Answer by jorel555(1290)

Mr. Pryme leaves at noon, driving 50 kmh, giving him a 50 km headstart. Thus:
50t+60t=226-50
110t=176
t=176/110 hours after 1:00 p.m.=96 minutes after 1:00 p.m.!!!!!!!!!!!!!!!

Answer by MathTherapy(10552)


At noon mr.Pryme decides to live kumasi for cape coast driving at 50km per hour.at 1:00pm Miss Anna live cape coast for kumasi driving at 60km per hour along the same route.If the two towns are 226km apart at what time will mr.Pryme meet Miss anna.
Time they met: